# Never-Ending Embodied Robot Learning

### v1

@@ Abstract
Relying on large language models (LLMs), embodied robots could perform complex multi-modal robot manipulation tasks from visual observations with powerful generalization ability. However, most visual behavior-cloning agents suffer from manipulation performance degradation and skill knowledge forgetting when adapting into a series of challenging unseen tasks. We here investigate the above challenge with NBCagent in embodied robots, a pioneering language-conditioned Never-ending Behavior-Cloning agent, which can continually learn observation knowledge of novel robot manipulation skills from skill-specific and skill-shared attributes. Specifically, we establish a skill-specific evolving planner to perform knowledge decoupling, which can continually embed novel skill-specific knowledge in our NBCagent agent from latent and low-rank space. Meanwhile, we propose a skill-shared semantics rendering module and a skill-shared representation distillation module to effectively transfer anti-forgetting skill-shared knowledge, further tackling catastrophic forgetting on old skills from semantics and representation aspects. Finally, we design a continual embodied robot manipulation benchmark, and several expensive experiments demonstrate the significant performance of our method. Visual results, code, and dataset are provided at: https://neragent.github.io.
Machine Learning, ICML
@@ 1 Introduction
Embodied robot learning (ERL) has attracted growing interests in merging machine learning with robot control system to solve various manipulation tasks. With the success of large language models, language-conditioned embodied robot can understand language instructions of users to achieve complex robot tasks and hold significant prospects for applications in industry, healthcare and home robot (Jiang et al., 2022; Brohan et al., 2022). Current language-conditioned behavior-cloning methods focus on address multi-modal data to efficiently execute complex manipulation tasks with visual observations. For instance, PerAct (Shridhar et al., 2023) utilizes a PerceiverIO Transformer (Jaegle et al., 2021) to encode language goals and RGB-D voxel observations, subsequently generating discretized robotic actions. As shown in Fig. 1, they could leverage robust scene understanding capabilities of large language models, and perform task generalization to achieve various novel manipulation tasks.
Figure 1: Illustration comparison between ERL and our NBCagent, where ERL performs novel task generalization over a fixed dataset, and our NBCagent can continually learn novel skill knowledge without catastrophic forgetting.
@@
However, the task generalization performance of most existing methods (e.g., PerAct (Shridhar et al., 2023)) is constrained when undertaking never-ending manipulation tasks and handling novel objects with intricate structures. Similar to the way humans acquire skills, this motivates us to enable robots to learn novel challenging manipulation skills in a continual learning manner. For instance, a home robot in the open-ended world is expected to consecutively learn various novel manipulation skills to meet the evolving needs of their owners. A trivial approach for this scenario involves enabling the robot to relearn these complex skills via visual observations, while avoiding forgetting previously acquired skills. However, the majority of existing methods assume training on a fixed dataset and only achieve this via storing old data and retraining model on all data, which leads to a large computational burden and high cost of memory storage, thereby being limited in real world.
@@
To address the aforementioned scenarios, we consider a practical challenging embodied robot learning problem, i.e., Never-ending Embodied Robot Learning (NERL), where embodied agent can perform continual learning on successive behavior-cloning manipulation skills and efficiently counteract catastrophic forgetting from old skills. Continual learning (Rebuffi et al., 2017; Yang et al., 2022; Sun et al., 2023) aims to continuously acquire knowledge from a successive data stream and has achieved remarkable performance in several computer vision tasks, such as image classification, object detection, image generation and so on. A naive solution for the NERL problem is to directly integrate continual learning and embodied robot learning together. However, different from learning category-wise knowledge focused by most existing methods, some attributes about skill-wise knowledge should be considered:
•  Skill-Specific Attribute originates from distinct manipulation sequences, object recognition, and scene understanding inherent in each unique skill. However, current continual learning methods neglect this attribute, which constrains their ability to continually learning novel skills.
•  Skill-Shared Attribute indicates that different skills possess shared knowledge. For instance, similar object recognition and scene understanding exist between different skills such as stacking bottles and opening bottle caps. Transferring skill-shared knowledge plays a key role in addressing skill-wise knowledge forgetting.
@@
To tackle the above-mentioned challenges, we propose a pioneering language-conditioned Never-ending Behavior-Cloning agent (i.e., NBCagent), which continually acquire skill-wise knowledge from skill-specific and skill-shared attributes with visual observations. To the best of our knowledge, this is an earlier attempt to explore never-ending embodied robot learning for multi-modal behavior-cloning robotic manipulation. To be specific, we propose a skill-specific evolving planner (SEP) to decouple the skill-wise knowledge in the latent and low-rank space, and focus on skill-specific knowledge learning. Additionally, we design a skill-shared semantics rendering module (SSR) and a skill-shared representation distillation module (SRD) to transfer skill-shared knowledge from semantics and representation aspects, respectively. Supervised by Neural Radiance Fields (NeRFs) and a vision foundation model, the SSR can complete skill-shared semantics in 3D voxel space across novel and old skills. Furthermore, the SRD can effectively distill knowledge between old and current models to align skill-shared representation. Several major contributions of our work are as follows:
@@
•  We take the earlier attempt to explore a novel real-world challenging problem called Never-ending Embodied Robot Learning (NERL), where we propose Never-ending Behavior-Cloning agent (i.e., NBCagent) to address the core challenges of skill-wise knowledge learning from skill-specific and skill-shared attributes.
•  We design a skill-specific evolving planner to decouple the skill-wise knowledge and continually embed skill-specific novel knowledge in our NBCagent. Moreover, a skill-shared semantic rendering module and a skill-shared representation distillation module is developed to learn skill-shared knowledge from semantics and representation aspects, respectively, and further overcome catastrophic forgetting.
•  We present a continual embodied robot manipulation benchmark for home robotic manipulation, which consists of two manipulation scenes, kitchen and living room. Qualitative experiments demonstrate the effectiveness and robustness of our proposed NBCagent.
Figure 2: Overview of the proposed NBCagent, where the perceiver model can continually learn novel manipulation skills. A skill-specific evolving planner is designed to learn skill-specific knowledge from latent and low-rank space. Meanwhile, we develop a skill-shared semantics rendering module and skill-shared representation distillation loss $\mathcal{L}_{srd}$ to transfer skill-shared knowledge from semantics and representation aspects.
@@ 2 Related Work
Robotic Manipulation. Recent works (Brohan et al., 2023; Chowdhery et al., 2023; Huang et al., 2022; Shah et al., 2023; Driess et al., 2023; Brohan et al., 2022; Zitkovich et al., 2023) have resulted in substantial advancements in the accomplishment of intricate tasks. VIMA (Jiang et al., 2022) proposes a novel multi-modal prompting scheme, which transforms a wide range of robotic manipulation tasks into a sequence modeling problem. Compared with directly using images as the manipulation input (Goyal et al., 2022; Shridhar et al., 2022), voxelizing 3D point clouds as a 3D representation (Shridhar et al., 2023; Ze et al., 2023; Goyal et al., 2023; James et al., 2022) can accomplish more complex tasks. PerAct (Shridhar et al., 2023) enables agent to perform better in robotic manipulation by voxelizing RGB-D images and discretizing output actions.
Continual Learning. Continual learning provides the foundation for the adaptive development of AI systems (Wang et al., 2023). The main approaches of continual learning can be categorized into three directions: Parameter regularization-based methods (Rebuffi et al., 2017; Li & Hoiem, 2017; Derakhshani et al., 2021; Douillard et al., 2020; Dong et al., 2023) balance the old and new tasks by adding more explicit regularization terms. Architecture-based methods (Jung et al., 2020; Wu et al., 2021; Wang et al., 2022; Toldo & Ozay, 2022) construct network parameters for different tasks. Replay-based methods include empirical replays (Bang et al., 2021; Rebuffi et al., 2017; Sun et al., 2022; Tiwari et al., 2022) and generative replays (Li et al., 2022; Xiang et al., 2019).
Some works focus on the improvement of robots by continual learning (Ayub & Wagner, 2023; Gao et al., 2021; Hafez & Wermter, 2023; Ayub & Fendley, 2022; Ayub et al., 2023). LOTUS (Wan et al., 2023) stores a few human demos of novel tasks into the growing skill library, enabling lifelong learning ability for robots. However, these methods cannot be applied to perform language-conditioned behaviour-cloning manipulation as they are incapable of handling multi-modal skill data. Furthermore, they neglect two inherent attributes in learning skill-wise knowledge, and suffer catastrophic forgetting on old skills.
@@ 3.1 Problem Definition and Overview
Problem Definition. Following traditional continual learning methods (Rebuffi et al., 2017; Douillard et al., 2020), we define a multi-modal skill data stream as $\mathcal{T}=\{\mathcal{T}^{m}\}_{m=1}^{M}$, where $M$ denotes the number of incremental tasks. Each incremental task consists of various robotic manipulation skills, resulting in a total of $N^{m}$ manipulation skills. The $m$-th incremental task $\mathcal{T}^{m}=\{\mathcal{D}^{m}_{i}\}_{i=1}^{N^{d}}$ consists of $N^{d}$ skill demonstrations. Specifically, each skill demonstration can be extracted to a set of keyframe actions, i.e., $\mathcal{D}^{m}_{i}=\{\mathbf{k}^{m,i}_{j}\}_{j=1}^{N^{k}},\mathbf{k}^{m,i}_{j}=\{\mathbf{a}^{m,i}_{j},\mathbf{r}^{m,i}_{j},\mathbf{l}^{m,i}\}$, where $N^{k}$ is the total keyframe action quantity. Given the current state in action space $\mathbf{a}$, the structured observation $\mathbf{r}$ and language instruction $\mathbf{l}$, the agent is expected to predict the next best keyframe action, which can be served as an action classification task (James et al., 2022). Additionally, the structured observation $\mathbf{r}$ is composed of the RGB-D images captured by a single front camera. An action state $\mathbf{a}$ can be divided into a discretized translation $\mathbf{a}_{tran}\in\mathbb{R}^{3}$, rotation $\mathbf{a}_{rot}\in\mathbb{R}^{(360/5)\times 3}$, gripper open state $\mathbf{a}_{grip}\in\{0,1\}$, and collision avoidance $\mathbf{a}_{col}\in\{0,1\}$. Here the rotation parameter $\mathbf{a}_{rot}$ entails the discretization of each rotation axis into a set of R = 5 bins. The collision avoidance parameter $\mathbf{a}_{col}$ provides guidance to the agent regarding the imperative need to avoid collisions.
@@
Overview. The overview of our NBCagent to learn skill-wise knowledge is shown in Fig. 2. When observing a novel incremental task $\mathcal{T}^{m}$, we initialize the perceiver model $\Theta^{m}_{p}$ for the current task utilizing the model $\Theta^{m-1}_{p}$ obtained from the last task and store $\Theta^{m-1}_{p}$ as a teacher model to perform our SRD module (we refer to PerceiverIO (Jaegle et al., 2021) as perceiver model for brevity). Following ER (Chaudhry et al., 2019), we build a memory buff $\mathcal{M}$ to replay only few samples form the previous tasks. In the $m$-th task, our NBCagent aims to execute all learned multi-modal robotic manipulation skills, achieved through iteratively optimizing the model on $\mathcal{T}^{m}$ and $\mathcal{M}$. Specifically, as shown in Fig. 2, given a RGB-D and language input, $\Theta^{m}_{p}$ first encode RGB-D input to obtain a deep 3D voxel utilizing a scaled-down voxel encoder $\mathcal{E}^{m}_{v}$. Then, we design a SSR module to transfer and complete skill-shared semantics across novel and old skills, which can effectively address catastrophic forgetting on old skills. After that, the patched voxel and language embeddings are sent to cross-attention blocks and self-attention blocks to perform feature extraction and semantics fusion, where we propose SEP to learn skill-specific knowledge from latent and low-rank space. Finally, we utilize a Q-function head to predict the state of the next keyframe in voxel space, where we develop a SRD loss $\mathcal{L}_{srd}$ to tackle catastrophic forgetting by aligning skill-shared representation.
3.2 Skill-Specific Evolving Planner
Aiming at learning category-wise knowledge, existing continual learning methods (Rebuffi et al., 2017; Douillard et al., 2020) assume that the knowledge acquired from novel tasks and that from previous tasks are mutually independent. Differently, for learning skill-wise knowledge, we consider decoupling the knowledge to the skill-shared knowledge and skill-specific knowledge. For instance, an embodied agent aims to stack the wine bottle after learning to open the wine bottle. It does not require relearning the skill-shared knowledge of scene understanding and object recognition; instead, the agent is expected to focus on acquiring the skill-specific knowledge related to the operation sequence of stacking and reducing forgetting on skill-shared knowledge. Considering this motivation, we design a skill-specific evolving planner (SEP) to perform knowledge decoupling, enabling effectively continual learning of novel skills. Specifically, we first develop an adaptive language semantic bank to retrieve the skill-specific language semantic information. In light of this information, SEP can encode the multi-modal input from skill-wise latent and low-rank space, and learn the novel knowledge through a skill-specific network.
First of all, when observing a novel skill, we utilize a language encoder $\mathcal{E}_{c}$ from CLIP (Radford et al., 2021) to encode the language instruction. This can obtain a skill-specific language semantic information $\mathbf{l}_{s}\in\mathbb{R}^{D^{c}}$, and text token information $\mathbf{l}_{x}\in\mathbb{R}^{N^{c}\times D^{x}}$, i.e., $\mathbf{l}_{s},\mathbf{l}_{x}=\mathcal{E}_{c}(\mathbf{l})$, where $D^{s}$ and $D^{x}$ represent the dimension of $\mathbf{l}_{s}$ and $\mathbf{l}_{x}$, and $N^{c}$ denotes the token length. Then, we compensate semantic information for our adaptive language semantic bank $\mathcal{B}$ during training via an exponential moving average strategy:
$\mathcal{B}[I,:]=(1-\mathcal{C}_{max})\mathcal{B}[I,:]+\mathcal{C}_{max}\mathbf{l}_{s},$ (1)
where $\mathcal{B}\in\mathbb{R}^{N^{b}\times D^{c}}$ is initialized by $N^{b}$ zero vectors. $\mathcal{C}\in\mathbb{R}^{N^{b}}$ is the cosine similarity matrix between the sentence information $\mathbf{l}_{s}$ and each vector in $\mathcal{B}$. if $\mathcal{C}_{max}>\delta$, we set $I=\mathrm{argmax}(\mathcal{C})$; otherwise, we set $I=\mathrm{nonzero}(\mathcal{B})+1$ and $\mathcal{C}_{max}=1$, where $\delta$ is a fixed threshold, and $\mathrm{nonzero}(\cdot)$ is a operation employed to calculate the number of nonzero vectors in $\mathcal{B}$. To this end, we can obtain the skill-wise code $I$ for the mini-batch training data, which plays a key role in performing skill-specific network training in the following.
Furthermore, different from existing multi-modal ERL methods (Shridhar et al., 2023; Ze et al., 2023) that only encode multi-modal input from a skill-shared latent space, our NBCagent considers to develop a dynamic skill-specific latent space $\mathbf{S}\in\mathbb{R}^{N^{s}\times N^{l}\times D^{s}}$, where $N^{s}$ denotes the number of learned skills, and $N^{l}$ represents the learnable latent vector quantity. NBCagent encodes these latent vectors with the multi-modal input utilizing a cross-attention layer to obtain the final feature. Specifically, following (Shridhar et al., 2023), we first to apply a scaled-down 3D convolution encoder to patch and encode a voxel observation $\mathbf{v}$ to obtain $\hat{\mathbf{v}}$, where $\mathbf{v}$ is obtained by voxelization process from $\mathbf{r}$. Then, we concatenate the encoded proprioception of agent in current state $\mathbf{a}$ and voxel observation $\hat{\mathbf{v}}$ to obtain $\mathbf{p}\in\mathbb{R}^{N^{p}\times D}$. In light of this, we employ a cross-attention layer to perform semantics interaction and obtain the cross-attention feature $\mathbf{F_{c}}\in\mathbb{R}^{(N^{p}+N^{c})\times D}$:
$\!\!\mathbf{F}_{c}=\rho(\frac{\mathrm{Cat}(\mathbf{p},\mathbf{c}_{x})\mathbf{W}_{q}(\mathbf{S}[I,:]\mathbf{W}_{k})^{\top}}{\sqrt{d}})(\mathbf{S}[I,:]\mathbf{W}_{v}),$ (2)
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}\in\mathbb{R}^{D\times D}$ are linear projection layers, and $\mathrm{Cat}(\cdot)$ denotes the concatenation operation. $\rho$ indices the softmax function and $d$ is a scaling factor. Then, we apply two additional linear projection layers $\mathbf{W}_{o}$ to handle the cross-attention feature $\mathbf{F}_{c}$. Similarly, we apply a series of self-attention blocks and a cross-attention decoder to further extract feature. In light of this, NBCagent can continually encode the novel multi-modal input from a skill-specific latent space, which is beneficial to learn some skill-specific knowledge from latent space.
Algorithm 1 Pipeline of Our ESP.
0:  language embedding $\mathbf{l}_{s}$;
0:  $\mathbf{S}[I,:],\mathbf{W}[I,:]$;
2:  if $\mathcal{C}_{max}>\delta$ then
3:     $I\xleftarrow{}\mathrm{argmax}(\mathcal{C});$
4:     Update $\mathcal{B}$ by Eq. (1);
5:     Return: $\mathbf{S}[I,:],\mathbf{W}[I,:]$;
6:  else
7:     $I\xleftarrow{}\mathrm{nonzero}(\mathcal{B}+1);$
8:     Expand $\mathcal{B}$ by Eq. (1);
9:     Randomly initialize $\mathbf{S}[I,:],\mathbf{W}[I,:]$;
10:     Return: $\mathbf{S}[I,:],\mathbf{W}[I,:]$;
11:  end if
Considering the limitation in representing skill-specific knowledge from latent space, we further explore to learn skill-specific knowledge from low-rank space. Specifically, we introduce a low-rank adaptation layer (LoRA) (Hu et al., 2021) that can learn skill-specific knowledge in an efficient manner. For $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$ in each attention block, we design a set of skill-specific LoRA layers to perform skill-specific forward and obtain the final output feature $\mathbf{F}_{x}$ as follows:
$\mathbf{F}_{x}=\mathbf{X}\mathbf{W}+\mathbf{X}\mathbf{W}_{r}[I,:]=\mathbf{X}\mathbf{W}+\mathbf{X}\mathbf{W}_{a}[I,:]\mathbf{W}_{b}[I,:],$ (3)
where $\mathbf{W}\in\mathbb{R}^{D\times D}$ represents one of $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v},\mathbf{W}_{o}$ and $\mathbf{X}\in\mathbb{R}^{(N^{v}+N^{p})\times D}$ denotes the input feature for these projection layers. $\mathbf{W}_{r}[I,:]=\mathbf{W}_{a}[I,:]\mathbf{W}_{b}[I,:]$ is a low-rank decomposition, where $\mathbf{W}_{a}\in\mathbb{R}^{N^{l}\times D\times N^{r}}$ is initialized in a random Gaussian manner, $\mathbf{W}_{b}\in\mathbb{R}^{N^{l}\times N^{r}\times D}$ is initialized by zero and $N^{r}$ is a hyper-parameter controlling the size of LoRA layers. On the one hand, obviously, $\mathbf{W}$ is shared among all skills and expected to learn skill-shared knowledge. On the other hand, each $\mathbf{W}_{r}$ is executed to learn different novel skills and perform skill-specific forward, resulting in continually embedding skill-specific knowledge to our NBCagent. Specifically, we summary the process of our SEP in Algorithm 1.
Algorithm 2 Optimization Pipeline of Our NBCagent.

### v2

# Never-Ending Behavior-Cloning Agent for Robotic Manipulation

@@ Abstract
Relying on multi-modal observations, embodied robots could perform multiple robotic manipulation tasks in unstructured real-world environments. However, most language-conditioned behavior-cloning agents still face existing long-standing challenges, i.e., 3D scene representation and human-level task learning, when adapting into new sequential tasks in practical scenarios. We here investigate these above challenges with NBAgent in embodied robots, a pioneering language-conditioned Never-ending Behavior-cloning Agent. It can continually learn observation knowledge of novel 3D scene semantics and robot manipulation skills from skill-shared and skill-specific attributes, respectively. Specifically, we propose a skill-shared semantic rendering module and a skill-shared representation distillation module to effectively learn 3D scene semantics from skill-shared attribute, further tackling 3D scene representation overlooking. Meanwhile, we establish a skill-specific evolving planner to perform manipulation knowledge decoupling, which can continually embed novel skill-specific knowledge like human from latent and low-rank space. Finally, we design a never-ending embodied robot manipulation benchmark, and expensive experiments demonstrate the significant performance of our method.
*Corresponding Author.
@@ 1 Introduction
Robot learning has attracted growing interests in integrating machine learning with robot control system to solve various robotic tasks, such as manipulation [39], navigation [36], mapping and localization [27]. In recent decades, behavior-cloning agents [17, 53] have demonstrated notable success in effective training with only a few demonstrations and direct implementation on real robots. Benefiting from vision-language observations, current multi-task behavior-cloning methods focus on utilizing multi-modal data to efficiently execute complex manipulation tasks with visual observations. For instance, PerAct [39] utilizes a PerceiverIO Transformer [21] to encode language goals and RGB-D voxel observations, subsequently generating discretized robotic actions.
@@
However, most existing language-conditioned behavior-cloning methods [21, 17, 26] assume that the skills mastered by a robot remain unchanged over time in real-world 2D scenarios, and focus on training policy model on a fixed set of manipulation tasks. They are resource-constrained when undertaking new skills and handling novel objects with intricate structures [44]. To address this scenario, the embodied robots—like human—should be capable of learning novel challenging manipulation skills in a continual learning manner. For instance, a home robot in the open-ended world is expected to consecutively learn various novel manipulation skills with demonstrations of behavior, and timely meet the evolving needs of its owners. Generally, a trivial approach for this scenario involves retraining the robot on all past and novel data, which leads to a large computational burden and high cost of memory storage, thereby is limited in real world unstructured 3D applications.
@@
To address the aforementioned scenarios, we consider a practical challenging robot learning problem, i.e., Never-ending Behavior-cloning Robot Learning (NBRL), where an agent can perform continual learning on successive behavior-cloning manipulation tasks and counteract catastrophic forgetting from learned skills. Different from learning category-wise knowledge focused by most existing methods [33, 44, 11], we here take the attempt to rethink some attributes about skill-wise knowledge:
•  Skill-Shared Attribute indicates that different complex skills possess shared knowledge, such as similar manipulated objects and 3D scenes understanding and so on. We primarily investigate the consistency of skill-shared knowledge across various skills on semantics of 3D scenes, which plays a key role in addressing skill forgetting and achieving comprehensive understanding on 3D scenes.
•  Skill-Specific Attribute originates from various factors such as distinct manipulation sequences, object recognition, motion primitives, and other distinct elements within each unique skill. Focusing the embodied robot to learn skill-specific knowledge can effectively perform novel skills learning and tackle forgetting on past learned manipulation skills or tasks.
@@
To tackle the above-mentioned challenges, we propose a pioneering language-conditioned Never-ending Behavior-cloning Agent (i.e., NBAgent), which can continually acquire skill-wise knowledge from both skill-shared and skill-specific attributes. To the best of our knowledge, this is an earlier attempt to explore continual learning for multi-modal behavior-cloning robotic manipulation. To be specific, we design a skill-shared semantic rendering module (SSR) and a skill-shared representation distillation module (SRD) to transfer skill-shared knowledge on semantics of 3D scenes. Supervised by Neural Radiance Fields (NeRFs) and a vision foundation model, SSR can transfer skill-shared semantic from 2D space into 3D space across novel and old skills; SRD can effectively distill skill-shared knowledge between old and current models to well align voxel representation. Additionally, we propose a skill-specific evolving planner (SEP) to decouple the skill-wise knowledge in the latent and low-rank space, and focus on skill-specific knowledge learning to alleviate manipulation skill forgetting. Several major contributions of our work are as follows:
@@
•  We take the earlier attempt to explore a practical challenging problem called Never-ending Behavior-cloning Robot Learning (NBRL), where we propose Never-ending Behavior-cloning Agent (i.e., NBAgent) to address the core challenges of skill-wise knowledge learning in 3D scene from skill-shared and skill-specific attributes.
•  We develop a skill-shared semantic rendering module and a skill-shared representation distillation module to capture the 3D semantic knowledge, which can transfer skill-shared semantic of 3D scenes to overcome 3D reasoning overlooking in continual learning.
•  We design a skill-specific evolving planner to perform skill-specific knowledge learning, which can decouple the skill-wise knowledge into latent and low-rank space, and continually embed novel skill-specific knowledge. Comprehensive experiments conducted on our proposed NBRL benchmark for home robotic manipulation demonstrate the effectiveness and robustness of our method.
@@ 2 Related Work
Robotic Manipulation. Recent works [6, 8, 20, 37, 14, 5, 54] have resulted in substantial advancements in the accomplishment of intricate tasks. VIMA [24] proposes a novel multi-modal prompting scheme, which transforms a wide range of robotic manipulation tasks into a sequence modeling problem. Compared with directly using images as the manipulation input [16, 38], voxelizing 3D point clouds as a 3D representation [39, 53, 17, 23] can accomplish more complex tasks. PerAct [39] enables agent to perform better in robotic manipulation by voxelizing RGB-D images and discretizing output actions. GNFactor [53] can generalize neural radiance field rendering through joint training, improving the model’s generalization ability.
Continual Learning. Continual learning provides the foundation for the adaptive development of AI systems [46, 10]. The main approaches of continual learning can be categorized into three directions: Parameter regularization-based methods [33, 30, 9, 12, 11] balance the old and new tasks by adding more explicit regularization terms. Architecture-based methods [25, 48, 45, 43] construct network parameters for different tasks. Replay-based methods include empirical replays [4, 33, 41, 42] and generative replays [29, 49, 40]. Some works focus on the improvement of robots by continual learning [3, 15, 18, 2, 1]. LOTUS [44] stores a few human demos of novel tasks into the growing skill library, enabling lifelong learning ability for robots. However, these methods are build on 2D vision observations, and cannot be applied to perform language-conditioned behaviour-cloning manipulation with multi-modal data.
@@ 3.1 Problem Definition and Overview
Problem Definition. By following traditional continual learning methods [33, 12], we define a multi-modal manipulation skill data stream as $\mathcal{T}=\{\mathcal{T}^{m}\}_{m=1}^{M}$, where $M$ denotes the number of continual skill learning tasks. The objective of NBAgent is to execute all acquired skills after observing $\mathcal{T}$. Specifically, each skill learning task consists of various robotic manipulation skills. The $m$-th continual skill learning task $\mathcal{T}^{m}=\{\mathcal{D}^{m}_{i}\}_{i=1}^{N^{d}}$ consists of $N^{d}$ skill demonstrations. Each skill demonstration $\mathcal{D}^{m}_{i}$ can be extracted to a set of keyframe actions [39], i.e., $\mathcal{D}^{m}_{i}=\{\mathbf{k}^{m,i}_{j}\}_{j=1}^{N^{k}},\mathbf{k}^{m,i}_{j}=\{\mathbf{a}^{m,i}_{j},\mathbf{r}^{m,i}_{j},\mathbf{l}^{m,i}\}$, where $N^{k}$ is the total keyframe action quantity. Given the current state in action space $\mathbf{a}$, the structured observation $\mathbf{r}$ and language instruction $\mathbf{l}$, the agent is expected to predict the next best keyframe action, which can be served as an action classification task [23]. Additionally, the structured observation $\mathbf{r}$ is composed of the RGB-D images captured by a single front camera. An action state $\mathbf{a}$ can be divided into a discretized translation $\mathbf{a}_{tran}\in\mathbb{R}^{3}$, rotation $\mathbf{a}_{rot}\in\mathbb{R}^{(360/5)\times 3}$, gripper open state $\mathbf{a}_{grip}\in\{0,1\}$, and collision avoidance $\mathbf{a}_{col}\in\{0,1\}$. Here the rotation parameter $\mathbf{a}_{rot}$ entails the discretization of each rotation axis into a set of R = 5 bins. The collision avoidance parameter $\mathbf{a}_{col}$ provides guidance to the agent regarding the imperative need to avoid collisions.
Figure 1: Overview of the proposed NBAgent. It consists of a skill-shared semantic rendering module and a skill-shared representation distillation loss $\mathcal{L}_{\mathrm{SRD}}$ to transfer skill-shared knowledge on semantics of 3D scenes, and a skill-specific evolving planner to learn skill-specific knowledge.
@@
Overview. The overview of NBAgent to learn skill-wise knowledge is shown in Fig. 1. When observing a novel sequential skill learning task $\mathcal{T}^{m}$, we initialize the perceiver model $\Theta^{m}_{p}$ for the current task utilizing the model $\Theta^{m-1}_{p}$ obtained from the last task, and store $\Theta^{m-1}_{p}$ as a teacher model to compute the SRD loss. Given a RGB-D image and language instruction, as shown in Fig. 1, $\Theta^{m}_{p}$ firstly encodes RGB-D input to obtain a deep 3D voxel via utilizing a scaled-down voxel encoder. We here design a SSR module to transfer skill-shared semantic of 3D scenes across novel and past encountered skills. Afterwards, the patched voxel and language embeddings are input to cross-attention blocks and self-attention blocks to perform feature extraction and semantics fusion, where we propose SEP to learn skill-specific knowledge from latent and low-rank space. Finally, we utilize a Q-function head $\mathbf{Q}^{m}$ to predict the state of the next keyframe in voxel space, where a SRD loss $\mathcal{L}_{\mathrm{SRD}}$ is developed to tackle catastrophic forgetting by aligning skill-shared voxel representation.
3.2 Skill-Shared Semantic Rendering Module
For language-conditional behaviour-cloning manipulation, a comprehensive semantics understanding of the 3D scenes [13] plays a key role in enabling agent to perform complicated manipulation skills. Especially in NBRL problem, which exists skill-shared semantic across various skills, such as 3D objects and scene semantics. The existence of overlooking on common semantic space makes these semantics incomplete, further resulting in catastrophic forgetting on past learned skills. Considering this motivation, we develop a skill-shared semantic rendering module (SSR) to transfer skill-shared semantic of 3D scenes, where a NeRF model and a vision foundation model are introduced to provide semantics supervision for this transfer process.
Concretely, we draw inspiration from 3D visual representation learning [53], and leverage a latent-conditioned NeRF architecture [52]. This architecture could synthesize RGB color $\mathbf{c}$ of a novel image views like traditional NeRF [31], and render the semantics $\mathbf{s}$ from the 3D voxel space as follows:
[…]
To distill skill-shared semantic in the $m$-th skill learning task, we initialize a NeRF model acquired from the last task and denote it as $\Theta_{n}^{m-1}$. Then, we feed the same input to obtain the pseudo ground truth $\hat{\mathbf{C}}$ by Eq. (2). We design a loss function to supervise the reconstruction process as follows:
$\mathcal{L}_{\mathrm{C}}=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\mathbf{Y}_{c}(\mathbf{r})\|^{2}_{2}+\beta\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{C}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{C}}(\mathbf{r},\mathbf{v}_{s})\|^{2}_{2}\cdot\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}},$ (3)
where $\mathbf{Y}_{c}$ indicates the ground truth color and $\mathcal{R}$ is the set of all camera rays. $\alpha$ is the hyper-parameter to control the weight of loss function. $\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}}$ is defined such that $\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}}=1$, when the condition $\mathbf{v}_{s}\notin\mathcal{T}^{m}$ is satisfied, and $\mathbb{I}_{\mathbf{v}_{s}\notin\mathcal{T}^{m}}=0$ otherwise.
Considering the insufficiency in capturing skill-shared semantic by reconstructing novel views, we introduce a pre-trained visual foundation model that contains robust scene semantics to provide supervision. Relying on being pre-trained on large-scale vision-language dataset, Stable Diffusion model [35] can possess robust intrinsic representational capabilities, which is consequently utilized for semantic representation in segmentation and classification tasks [50, 28]. In light of this, we employ Stable Diffusion model $\Theta_{u}$ to extract vision-language semantics for supervision. Given a input view, i.e., $\mathbf{Y}_{c}$, we perform a one-step noise adding process to obtain a noisy image $\mathbf{Y}_{c,t}$. Then we utilize diffusion model $\Theta_{u}$ to collect vision-language semantic feature as the ground truth $\hat{\mathbf{F}}_{s}$:
$\mathbf{Y}_{c,t}(\mathbf{r}):=\sqrt{\alpha_{t}}\mathcal{E}_{v}(\mathbf{Y}_{c}(\mathbf{r}))+\sqrt{1-\alpha_{t}}\epsilon,~{}~{}\hat{\mathbf{F}}_{s}(\mathbf{r},\mathbf{l}_{p})=\Theta_{u}(\mathbf{Y}_{c,t}(\mathbf{r}),\mathcal{E}_{c}(\mathbf{l}_{p})),$ (4)
where $\mathcal{E}_{v}$ is a VAE encoder to encode image $\mathbf{Y}_{c}$ from pixel space to latent semantic space. $t$ represents the diffusion process step, $\epsilon\sim\mathcal{N}(0,1)$ and $\alpha_{t}$ is designed to control the noise schedule. $\mathbf{l}_{p}$ denotes the language prompt modified from task description $\mathbf{l}$. To perform skill-shared semantic transfer, we align the rendered semantic feature $\mathbf{F}_{s}$ and diffusion feature $\hat{\mathbf{F}}_{s}$, and obtain the major objective of our SSR module as follows:
$\mathcal{L}_{\mathrm{SSR}}=\mathcal{L}_{\mathrm{C}}+\lambda_{1}\mathcal{L}_{\mathrm{S}},~{}~{}\mathcal{L}_{\mathrm{S}}=\underset{\mathbf{r}\in\mathcal{R}}{\sum}\|\mathbf{F}_{s}(\mathbf{r},\mathbf{v}_{s})-\hat{\mathbf{F}}_{s}(\mathbf{r},\mathbf{l}_{p})\|^{2}_{2}.$ (5)
Algorithm 1 Optimization Pipeline of NBAgent.